# Data-Efficient Quantum Noise Modeling via Machine Learning

### v1

@@ II.2.3 Readout errors
[…]
@@ II.3 Parameter optimization
The error model introduced in the previous section contains 20 free parameters $\bm{\theta}$. Our goal is to choose the parameters of the extended model such that the model accurately reproduces sample distributions $P_{i}$ obtained from executing circuits $C_{i}\in\mathcal{C}$ on a specific quantum computing backend. This can be formalized as identifying the parameter configuration $\bm{\theta}^{*}$ that minimizes the Hellinger distance [32] between simulation results and the results obtained on a real device. The Hellinger distance $D_{H}$ ranges from 0 to 1, with 0 indicating identical distributions, and is defined as
@@
$\bm{\theta}^{*}=\text{argmin}_{\bm{\theta}}\left\{\frac{1}{|{\mathcal{C}}|}\sum_{i}H[P_{i},Q_{i}(\bm{\theta})]\right\}\,.$ (15)
Here, $|{\mathcal{C}}|$ is the size of the training set, $Q_{i}(\bm{\theta})$ is the empirical distribution obtained by simulating the circuit with a noise model with parameters $\bm{\theta}$. Since the configuration space of the model is large and evaluating the Hellinger distance requires running noisy simulations, identifying $\bm{\theta}^{*}$ is not tractable using brute force search. We thus employ Bayesian optimization (BO) to efficiently obtain estimates of $\bm{\theta}^{*}$. Bayesian optimization is a global optimization technique particularly effective for expensive-to-evaluate black-box objective functions [36]. It employs a probabilistic surrogate model to intelligently select parameter evaluations by maximizing an acquisition function that balances exploration of uncertain regions with exploitation of promising areas.
As a surrogate model, we use a tree-structured Parzen estimator [37], which uses adaptive, non-parametric Parzen window density models to distinguish promising from poor regions of the search space based on past evaluations. Then samples new configurations from the most promising areas to enable efficient hyperparameter optimization in complex, hierarchical domains. We benchmark the BO against random search (RS) in which parameter points are randomly selected from the search space to provide unbiased model estimates. Both methods are implemented using Optuna [38].
@@ III Experimental Results
We benchmark our method using data collected from multiple IBM QPUs by executing quantum algorithms with varying qubit counts. The chosen systems ibmq_kolkata, ibmq_mumbai, and ibmq_ehningen comprise 27 qubits and use CNOT gate as their native two-qubit entangling operation, whereas ibm_kingston features 156 qubits and uses CZ gate. The benchmarks include the QAOA and VQE, which are critical classes of variational algorithms designed for near-term hardware. We further include random circuits to evaluate the framework’s suitability for unstructured quantum computations. Moreover, we partition our benchmark circuits by size into a small-size training set (4–6 qubits) and a large-size validation set (7–9 qubits) for testing the model’s ability to extrapolate from smaller systems to predict the behavior of larger, more complex systems.
@@ III.1 QAOA
where $Z_{i}$ are Pauli-$Z$ operators and the coefficients $c_{i},c_{ij}$ reflect the problem structure. In this work, we employ QAOA applied to portfolio optimization problems, which involves fully connected two-qubit gates; further details can be found in [15, 16, 39, 4].
For each QAOA circuit, we use the constrained optimization by linear approximation (COBYLA) [40] to choose angles $\bm{\gamma}$ and $\bm{\beta}$. Additionally, we use SWAP networks [15, 41] on linear topologies to address the connectivity constraints of the QPU, which enables all compiled circuits to share the same fundamental gate structure and qubit connectivity pattern, regardless of the target device. This standardization allows us to control for variability in circuit design and compilation strategies.
We quantify circuit depth and gate counts of benchmark circuits executed on ibmq_kolkata, ibmq_mumbai, and ibmq_ehningen to assess their complexity and noise sensitivity. As detailed in Table 2, the training set comprises circuits with 4–6 qubits and depths of $27.0\pm 3.4$, while the prediction set exhibits 7–9 qubits and depths of $39.0\pm 3.4$. The scaling of CNOT gates, from $27.0\pm 10.7$ in training to $78.0\pm 18.4$ in prediction, reflects the increased entanglement complexity and noise accumulation in larger circuits.
@@ III.1.1 Single-device characterization
We evaluate the method on ibmq_kolkata. On the training set, our parameterized noise model optimized with BO and RS achieves an average Hellinger distance of 0.095 and 0.118, respectively. This represents a significant reduction in predictive error of 45.92% and 32.50% compared to the default model’s distance of 0.175, as shown in Fig. 2(a). Crucially, the model’s performance generalizes robustly to the validation set (7–9 qubits) (Fig. 2(b)). For these larger circuits, our model achieves an average Hellinger distance of 0.135 (BO) and 0.148 (RS), corresponding to a reduction of 50.91% and 46.22% over the default model’s distance of 0.275. These results also show that BO consistently outperforms RS in both the training and prediction processes, demonstrating its high optimization efficiency.
@@ III.1.2 Cross-platform comparative analysis
where $\epsilon_{i}$ are gate and readout errors. The increase in complexity from training set to prediction set corresponds to a significant circuit fidelity degradation from about $78\%$ to about $50\%$, as shown in Table 3, underscoring the challenge of extrapolating noise models to larger circuit regimes where errors become more pronounced. Notably, the physical qubit allocation, driven by device-specific transpilation strategies, reveals that prediction circuits span a broader device region compared to training circuits, potentially introducing greater variability from spatial heterogeneity in device error rates. We observe that ibmq_ehningen, compared to other QPUs, maps circuits across a significantly wider physical qubit topology across four different benchmark executions. The increased spatial complexity is probably responsible for the more modest performance gains. A potential reason is that our amount of training data is insufficient to fully characterize the more complex error landscape of an expanded qubit ensemble. Nevertheless, the systematic increase in circuit complexity from training to prediction provides a rigorous test of our noise model’s ability to extrapolate beyond its training regime. The key aspect of this extrapolation is the model’s capacity to characterize qubits absent from the training set. This is enabled because our framework avoids qubit-specific error rate mappings. Instead, it constructs a generative model of the device’s noise, optimizing global hyperparameters $\bm{\theta}$ that parameterize statistical distributions for gate errors from which individual qubit and gate parameters are drawn.
@@ III.1.3 Training efficiency study
This section examines the interplay between model performance and training dataset size to identify the minimal data volume needed for target accuracy with reduced resource demands. To assess our framework’s data requirements, we train and validate our noise model using four distinct QAOA datasets of varying sizes (9, 12, 15, and 18 circuits). The results reveal a remarkable degree of robustness and data efficiency. Across 9 to 18 distinct QAOA circuit sizes, circuit properties remain remarkably stable. The preserved complexity distribution across all dataset sizes, combined with consistent fidelity estimates, validates that our smallest 9-circuit training set adequately captures the noise characteristics necessary for model optimization; more details can be found in Table 6 in Appendix A. This structural invariance explains the robust performance shown in Fig. 5, where the framework achieves a consistently high reduction in $D_{H}$ of approximately 45% to 50% across all tested datasets, despite a $2\times$ difference in training data volume. Moreover, the optimization convergence profiles for 12, 15, and 18 circuit datasets are nearly indistinguishable, reaching the same optimal $D_{H}$ value at a similar rate, highlighting stability (see Fig. 6 in Appendix A). This key finding confirms that our method does not require extensive, experimentally costly training data to construct a high-fidelity noise model, highlighting its practical efficiency.
@@ III.2 VQE
To verify the efficacy of the method beyond QAOA, we perform a similar optimization for VQE circuits. We consider the task of minimizing the ground state energy of the cluster Hamiltonian
@@
where $\alpha_{i}$ and $\beta_{i}$ are real coefficients. In the following, we choose $\alpha_{k}=1/n$ ($\beta_{k}=1/n$) for all $k$ for $n=4,\dots,9$. Using a simple, hardware-efficient ansatz corresponding to the EfficientSU2 circuit in Qiskit, we optimize the parameters using the sequential least squares quadratic programming (SLSQP) [42] in a noiseless simulation and verify that the optimized circuits converge to the true ground state energy. We then execute the optimized circuits on the 156-qubit Heron R2 QPU ibm_kingston. For the execution, we group the circuits for each $n$ into two qubit-wise commuting groups and sample from each group with $30.000$ shots. This results in a total of $12$ circuits which we partition into six circuits ($n=4,5,6$) for training and six circuits for validation ($n=7,8,9$).
Performing three independent optimizations using $200$ iterations each, we obtain optimized parameters which result in an improvement of $43.99\%\pm 0.14\%$ in model fidelity for the training set and $18.58\%\pm 0.90\%$ for the validation set (up to 27% for a circuit prediction, see Fig. 7(a) in Appendix A). While the parameterized model aligns more closely with experiment, especially on the training circuits, the improvement on the validation set is less pronounced than in the QAOA case. This likely reflects differences between the devices from which the data were sampled. As detailed in Table 4, VQE circuits contain significantly fewer entangling gates (CZs) than QAOA circuits (CNOTs). Moreover, the ibm_kingston processor itself exhibits lower error rates. These factors contribute to inherently higher baseline circuit fidelities, resulting in substantially smaller Hellinger distances between the baseline model and experiment. On the validation set, the default model’s Hellinger distance is roughly half of that observed for the comparable model on, e.g., ibmq_kolkata. Consequently, the margins of improvement over the baseline are smaller. Nevertheless, these results show that the approach in this work is robust across device generations, qubit counts, and native gate sets, establishing a clear path forward for modeling noise in more complex VQE circuits.

### v2

@@ II.2.3 Readout errors
We extend the standard model of independent single-qubit readout errors to include a scalable approximation of correlated errors. In superconducting processors, readout correlations often arise from shared resonators or multiplexed electronics, potentially linking multiple qubits. However, modeling the full joint readout distribution of $n$ qubits requires a confusion matrix scaling as $2^{n}\times 2^{n}$. This exponential scaling is incompatible with our objective of a data-efficient, circuit-size-independent framework. Instead, we model correlations at the pairwise level as the leading-order correction to the independent model. We approximate the global readout confusion matrix $\mathcal{M}_{\text{global}}$ as a tensor product of independent single-qubit matrices $M^{(1)}$ and disjoint two-qubit matrices $M^{(2)}$:
$\mathcal{M}_{\text{global}}\approx\left(\bigotimes_{(q_{i},q_{j})\in\mathcal{S}}M_{ij}^{(2)}\right)\otimes\left(\bigotimes_{q_{k}\notin\mathcal{S}}M_{k}^{(1)}\right),$ (19)
where $\mathcal{S}$ is a set of disjoint qubit pairs. To construct $\mathcal{S}$, we employ a deterministic greedy maximal matching algorithm on the device coupling graph. This algorithm iterates through the edges of the coupling map (sorted by qubit index to ensure reproducibility) and selects a pair $(q_{i},q_{j})$ if neither qubit has been previously selected. This approach maximizes the coverage of correlated errors, prioritizing physically connected neighbors where crosstalk is typically strongest, while ensuring each qubit participates in at most one correlated channel.
For a selected pair $(q_{0},q_{1})$, we begin with the $4\times 4$ confusion matrix representing independent errors, $M_{\text{ind}}=M_{q_{1}}\otimes M_{q_{0}}$. We then construct the correlated confusion matrix $M_{ij}^{(2)}$ by injecting correlations into this product. Specifically, we model the $|00\rangle\leftrightarrow|11\rangle$ correlation, corresponding to a simultaneous correlated readout misassignment on both qubits. This is achieved by transferring a probability mass $\lambda_{\text{corr}}^{00\leftrightarrow 11}$ from the diagonal elements to the anti-diagonal elements (correlated error) within the subspaces of the affected states, such as
@@ II.2.4 Model parameterization and calibration integration
Crucially, our framework employs a hybrid parameterization strategy that integrates global learnable parameters with local calibration data. Instead of learning independent error rates for every qubit, which would result in an unscalable parameter space, the 20 optimization variables $\bm{\theta}$ function as global modifiers applied to the spatially varying, qubit-specific calibration inputs. The error channels leverage standard device calibration parameters, including relaxation times $T_{1}$, dephasing times $T_{2}$, gate durations and error rates of basis gates, which are sourced directly from routine vendor-provided characterization data. Utilizing these existing data requires no dedicated calibration circuits or additional QPU time from the user. This architecture ensures that the model naturally captures spatial heterogeneity. For instance, in the amplitude damping channel (Eq. 8), the relaxation time $T_{1}$ varies locally per qubit, while the offset $b_{\text{amp}}$ is shared globally. Consequently, for a fixed gate duration $t_{g}$, a qubit with lower $T_{1}$ is assigned a higher damping probability than one with higher $T_{1}$ despite identical $b_{\text{amp}}$. Learning global corrections to local physical priors avoids the prohibitive complexity of learning unique parameter sets for every qubit and qubit pair, while successfully modeling the non-uniform error landscape.
@@ II.3 Parameter optimization
The error model introduced in the previous section contains 20 free parameters $\bm{\theta}$. Our goal is to choose the parameters of the extended model such that the model accurately reproduces sample distributions $P_{i}$ obtained from executing circuits $C_{i}\in\mathcal{C}$ on a specific quantum computing backend. This can be formalized as identifying the parameter configuration $\bm{\theta}^{*}$ that minimizes the Hellinger distance [hellinger1909neue] between simulation results and the results obtained on a real device. The Hellinger distance $D_{H}$ is a bounded, symmetric metric for comparing discrete probability distributions and is numerically stable for finite-shot, sparse bitstring histograms. It ranges from 0 to 1, with 0 indicating identical distributions, and is defined as
@@
$\bm{\theta}^{*}=\text{argmin}_{\bm{\theta}}\left\{\frac{1}{n_{\rm tr}}\sum_{i=1}^{n_{\rm tr}}D_{H}[P_{i},Q_{i}(\bm{\theta})]\right\}\,.$ (23)
Here, ${n_{\rm tr}}=|{\mathcal{C}}|$ is the size of the training set, $Q_{i}(\bm{\theta})$ is the empirical distribution obtained by simulating the circuit with a noise model with parameters $\bm{\theta}$. We emphasize that Eq. (23) defines a global fit, meaning that a single parameter vector $\bm{\theta}$ is identified jointly over all circuits in $\mathcal{C}$, with no circuit-specific parameters. This structure enforces that $\bm{\theta}$ captures systematic, context-consistent error mechanisms shared by the algorithm family, rather than memorizing individual circuit outcomes, which potentially mitigates overfitting.
Since the configuration space of the model is large and evaluating the Hellinger distance requires running noisy simulations, identifying $\bm{\theta}^{*}$ is not tractable using brute force search. Moreover, gradient-based optimization faces practical limitations in this setting. Efficient gradient-based optimization via automatic differentiation would require end-to-end differentiability of the noisy-simulation stack with respect to the noise parameters. In our workflow based on standard simulators (e.g., Qiskit Aer), such gradients are not natively available; obtaining them would require substantial reimplementation or numerical gradient estimation via finite differences, incurring $\mathcal{O}(d)$ additional objective evaluations per update. With $d=20$ parameters in our model, this imposes a substantial computational overhead given that the noisy simulation itself is the primary bottleneck. Additionally, the noise landscape is often non-convex and can be multimodal due to the complex interplay between coherent and incoherent error channels, making gradient-based methods susceptible to local minima or sensitive to initialization and hyperparameters. We thus employ Bayesian optimization to efficiently obtain estimates of $\bm{\theta}^{*}$. BO is a machine learning-based strategy well-suited for expensive-to-evaluate black-box objective functions [ArchettiCandelieri2019]. Unlike standard gradient-based local search, which determines step directions based on local derivatives, BO employs a probabilistic surrogate model to intelligently select parameter evaluations by maximizing an acquisition function that balances exploration of uncertain regions with exploitation of promising areas.
As a surrogate model, we use a tree-structured Parzen estimator [bergstra2011algorithms], which uses adaptive, non-parametric Parzen window density models to distinguish promising from poor regions of the search space based on past evaluations. Then samples new configurations from the most promising areas to enable efficient hyperparameter optimization in complex, hierarchical domains. We benchmark the BO against random search (RS) in which parameter points are randomly selected from the search space to provide unbiased model estimates. Both methods are implemented using Optuna [optuna_2019].
@@ III Experimental Results
We benchmark our method using data collected from multiple IBM QPUs by executing quantum algorithms with varying qubit counts. The chosen systems ibmq_kolkata, ibmq_mumbai, and ibmq_ehningen comprise 27 qubits with CNOT as the basis gate, physically implemented via echoed cross-resonance and single-qubit rotations. Notably, ibmq_ehningen supports two distinct physical implementations of the CNOT gate [ji2023optimizing]. In contrast, ibm_kingston features 156 qubits and utilizes native CZ gates implemented via tunable couplers. Our noise model characterizes these basis gates, as all circuits must be decomposed into the basis gate set for execution. The benchmarks include the QAOA and VQE, which are critical classes of variational algorithms designed for near-term hardware. We further include random circuits to evaluate the framework’s suitability for unstructured quantum computations. Moreover, we partition our benchmark circuits by size into a small-size training set (4–6 qubits) and a large-size validation set (7–9 qubits) for testing the model’s ability to extrapolate from smaller systems to predict the behavior of larger, more complex systems.
@@ III.1 QAOA
where $Z_{i}$ are Pauli-$Z$ operators and the coefficients $c_{i},c_{ij}$ reflect the problem structure. In this work, we employ QAOA applied to portfolio optimization problems, which involves fully connected two-qubit gates; further details can be found in [ji2025algorithm, ji2023improving, ji2023optimizing, brandhofer2023benchmarking].
For each QAOA circuit, we use the constrained optimization by linear approximation (COBYLA) [powell1994direct] to choose angles $\bm{\gamma}$ and $\bm{\beta}$. Additionally, we use SWAP networks [ji2025algorithm, montanez2025optimizing, weidenfeller2022scaling] on linear topologies to address the connectivity constraints of the QPU, which enables all compiled circuits to share the same fundamental gate structure and qubit connectivity pattern, regardless of the target device. This standardization allows us to control for variability in circuit design and compilation strategies.
We quantify circuit depth and gate counts of benchmark circuits executed on ibmq_kolkata, ibmq_mumbai, and ibmq_ehningen to assess their complexity and noise sensitivity. As detailed in Table 2, the training set comprises circuits with 4–6 qubits and depths of $27.0\pm 3.4$, while the prediction set exhibits 7–9 qubits and depths of $39.0\pm 3.4$. The scaling of CNOT gates, from $27.0\pm 10.7$ in training to $78.0\pm 18.4$ in prediction, reflects the increased entanglement complexity and noise accumulation in larger circuits. While the increase in circuit width (4–6 to 7–9 qubits) is moderate, this nearly $3\times$ increase in entangling gates drives the validation circuits into a significantly lower circuit fidelity regime (see Table 3), ensuring that the model is rigorously tested on its ability to generalize error dynamics into a more noise-affected operating regime distinct from the high circuit fidelity training regime.
@@ III.1.1 Single-device characterization
We evaluate the method on ibmq_kolkata. On the training set, our parameterized noise model optimized with BO and RS achieves an average Hellinger distance of 0.095 and 0.118, respectively. This represents a significant reduction in predictive error of 45.71% and 32.57% compared to the default model’s distance of 0.175, as shown in Fig. 2(a). Crucially, the model’s performance generalizes robustly to the validation set (7–9 qubits) (Fig. 2(b)). For these larger circuits, our model achieves an average Hellinger distance of 0.135 (BO) and 0.148 (RS), corresponding to a reduction of 50.91% and 46.18% over the default model’s distance of 0.275. These results also show that BO consistently outperforms RS in both the training and prediction processes, demonstrating its high optimization efficiency.
@@ III.1.2 Cross-platform comparative analysis
where $\epsilon_{i}$ are gate and readout errors. The increase in complexity from training set to prediction set corresponds to a significant circuit fidelity degradation from about $78\%$ to about $50\%$, as shown in Table 3, underscoring the challenge of extrapolating noise models to larger circuit regimes where errors become more pronounced. Notably, the physical qubit allocation, driven by device-specific transpilation strategies, reveals that prediction circuits span a broader device region compared to training circuits, potentially introducing greater variability from spatial heterogeneity in device error rates. We observe that ibmq_ehningen, compared to other QPUs, maps circuits across a significantly wider physical qubit topology across four different benchmark executions. The increased spatial complexity is probably responsible for the more modest performance gains. A potential reason is that our amount of training data is insufficient to fully characterize the more complex error landscape of an expanded qubit ensemble. Nevertheless, the systematic increase in circuit complexity from training to prediction provides a meaningful test of our noise model’s ability to extrapolate beyond its training regime. This key extrapolation is enabled because our framework avoids qubit-specific error rate mappings. Instead, it constructs a generative model of the device’s noise, optimizing global hyperparameters $\bm{\theta}$ that parameterize statistical distributions for gate errors from which individual qubit and gate parameters are drawn.
In our experiments, we maintain circuit fidelities $F\gtrsim 50\%$ in both the training phase (4–6 qubits) and the larger-scale prediction phase (7–9 qubits), so their output distributions remain far from uniform and retain structure that is sensitive to coherent and correlated error mechanisms. By contrast, circuits with substantially larger width (e.g., 20–30 qubits) at comparable depth on the same devices are expected to operate close to the noise floor, yielding nearly uniform outcome distributions that obscure detailed differences between noise models and strongly reduce the discriminative power of metrics such as the Hellinger distance. Our conclusions should therefore be understood as applying to this intermediate-fidelity regime; extending the analysis to significantly larger mapped circuits will require dedicated experiments and is left for future work.
@@ III.1.3 Training efficiency study
This section examines the interplay between model performance and training dataset size to identify the minimal data volume needed for target accuracy with reduced resource demands. To assess our framework’s data requirements, we train and validate our noise model using four distinct QAOA datasets of varying sizes (9, 12, 15, and 18 circuits). The results reveal strong robustness and data efficiency. Across training-set sizes from 9 to 18 circuits, circuit properties remain stable. The preserved complexity distribution across all dataset sizes, combined with consistent fidelity estimates, suggests that our smallest 9-circuit training set sufficiently constrains the model for this benchmark; more details can be found in Table 6 in Appendix A. This structural invariance explains the robust performance shown in Fig. 5, where the framework achieves a consistently similar relative reduction in the mean $D_{H}$ of approximately 45% to 50% across all tested datasets, despite a $2\times$ difference in training data volume. Moreover, the optimization convergence profiles for 12, 15, and 18 circuit datasets are nearly indistinguishable, reaching the same optimal $D_{H}$ value at a similar rate, highlighting stability (see Fig. 6 in Appendix A).
These results demonstrate that the framework is data-efficient and provides empirical evidence of performance saturation with respect to training-set size, with no indication of overfitting in the tested regime. Within the tested range (9–18 training circuits), performance stabilizes: increasing the training set from 9 to 18 circuits yields no appreciable change in the optimal $D_{H}$ and produces nearly indistinguishable convergence behavior (Fig. 6). This is consistent with adequate constraint diversity: circuits with varying depths and two-qubit-gate densities probe distinct error channels, providing complementary constraints on the shared 20-parameter model, while physics-informed functional forms restrict effective capacity by confining $\bm{\theta}$ to physically realistic mechanisms. Out-of-distribution validation, where improvements match or in several cases exceed training across devices, further confirms robust generalization.
@@ III.2 VQE
To verify the efficacy of the method beyond QAOA, we perform a similar optimization for VQE circuits on ibm_kingston, a 156-qubit Heron R2 processor featuring IBM’s heavy-hex topology. We consider the task of minimizing the ground state energy of the cluster Hamiltonian
@@
where $\alpha_{i}$ and $\beta_{i}$ are real coefficients. In the following, we choose $\alpha_{k}=1/n$ ($\beta_{k}=1/n$) for all $k$ for $n=4,\dots,9$. Using a simple, hardware-efficient ansatz corresponding to the EfficientSU2 circuit in Qiskit, we optimize the parameters using the sequential least squares quadratic programming (SLSQP) [kraft1988software] in a noiseless simulation and verify that the optimized circuits converge to the true ground state energy. We then execute the optimized circuits on the 156-qubit Heron R2 QPU ibm_kingston. For the execution, we group the circuits for each $n$ into two qubit-wise commuting groups and sample from each group with $30.000$ shots. This results in a total of $12$ circuits which we partition into six circuits ($n=4,5,6$) for training and six circuits for validation ($n=7,8,9$).
As in the QAOA study, we apply the same parameterized noise model architecture to the VQE experiments on ibm_kingston; however, the parameters $\bm{\theta}$ are learned solely from the VQE training data specific to ibm_kingston. Performing three independent optimizations using $200$ iterations each, we obtain optimized parameters which result in an improvement of $43.99\%\pm 0.14\%$ in model fidelity for the training set and $18.58\%\pm 0.90\%$ for the validation set (up to 27% for a circuit prediction, see Fig. 7(a) in Appendix A). While the parameterized model aligns more closely with experiment, especially on the training circuits, the improvement on the validation set is less pronounced than in the QAOA case. This likely reflects differences between the devices from which the data were sampled. As detailed in Table 4, VQE circuits contain significantly fewer entangling gates (CZs) than QAOA circuits (CNOTs). Moreover, the ibm_kingston processor itself exhibits lower error rates. These factors contribute to inherently higher baseline circuit fidelities, resulting in substantially smaller Hellinger distances between the baseline model and experiment. On the validation set, the default model’s Hellinger distance is less than half of that observed for the comparable model on the older generation devices used for QAOA. Consequently, the margins of improvement over the baseline are smaller. Nevertheless, these results demonstrate that our framework is robust across diverse hardware architectures, accommodating distinct qubit counts and native gate sets, establishing a clear path forward for modeling noise in more complex VQE circuits.